OpenStudy (aquilla):

Please help with this! four fifth x minus three and three fourths x equals two and one half

OpenStudy (anonymous):

Alright first thing i would do on this is make your fractions match ...lets forget about x for a moment .... you have 4 /5 and 3 3/4 which as a mixed fraction is 9/4 correct?

OpenStudy (aquilla):

I think so

OpenStudy (anonymous):

well make 3 3/4 just a straight fraction and what do you get 3 x 4 bc is 3 4/4s and then an extra 3 which makes it 15/4 excuse me

OpenStudy (aquilla):

Okay

OpenStudy (aquilla):

uh, is that the answer?

OpenStudy (anonymous):

no its not..you are trying to find x...sorry my pc turned off

OpenStudy (anonymous):

it says equal 2 and 1/2 correct?

OpenStudy (aquilla):

yeah

OpenStudy (anonymous):

make that into just a fraction

OpenStudy (aquilla):

okay

OpenStudy (anonymous):

what did you get?

OpenStudy (aquilla):

5/2 Is this right?

OpenStudy (anonymous):

yes that is correct...the next thing you need to do is give all 3 fractions a common denominator ...4 5 and 2 have what number in common?

OpenStudy (aquilla):

20

OpenStudy (anonymous):

good ...now to turn them into correct fractions with that number as a denominator ..you multiply the numerator or top by the number you need to make the denominator 20 ..for example ..with 5/2 ...you need the number 10 to make the denominator a 20 correct? ..so you multiply the numerator and get 50/20

OpenStudy (aquilla):

Okay

OpenStudy (aquilla):

45/20, 75/20/, 50/20

OpenStudy (anonymous):

the first one i don't believe is correct

OpenStudy (aquilla):

16/20 sorry

OpenStudy (anonymous):

yep. now basically still don't put the x in ..but combine like terms..show me the equation without x in it and using those fractions

OpenStudy (aquilla):

16/20 - 75/20= 50/20?

OpenStudy (anonymous):

yes ...now combine like terms

OpenStudy (aquilla):

-53/20= 50/20?

OpenStudy (anonymous):

that would be -59/20 but yes

OpenStudy (aquilla):

Oops, sorry. I meant to put that.

OpenStudy (anonymous):

alright so now you have -59/20x = 50/20

OpenStudy (aquilla):

50/-59?

OpenStudy (anonymous):

(50/20)/ -(59/20) = x

OpenStudy (aquilla):

okay, thank you so much! I appreeciate it!

OpenStudy (anonymous):

no problem

OpenStudy (aquilla):

can I ask you another question please?

OpenStudy (anonymous):

yeah make a new qustion though and close this one

OpenStudy (aquilla):

k

OpenStudy (aquilla):

wait, I think I got it so far. I'll let you know if I need any help again. Thanks